OpenStudy (gorica):

how to expess y from the equation x+e^y=ln(x+y^2) ?

OpenStudy (solomonzelman):

No, Idk, I tried to do some staff, but I wonder how would this problem be done, sorry . I can tag someone though: @shamil98

OpenStudy (gorica):

if it's easier, you could express x instead of y

OpenStudy (shamil98):

sorry, i'm doing my own homework atm, got tons -.-

OpenStudy (solomonzelman):

Ok, @helper_edwin pliz...

OpenStudy (solomonzelman):

@helder_edwin (sorry)

OpenStudy (helder_edwin):

I tried it! There is no way to solve that equation neither for x nor for y.

OpenStudy (gorica):

Is there a way to find a derivative of that function? That's what I need in fact... I was trying to express y via x or oposite and than to find a derivative... But it seems it's impossible to do it like that :)

OpenStudy (helder_edwin):

in that case, there is no need for that. Just differentiate implicitly. do u know how to?

OpenStudy (helder_edwin):

Implicit differentiation is used precisely with functions for which it is imposible to express y in terms of x.

OpenStudy (gorica):

x'+(e^y)y'=(x+y^2)'/(x+y^2)?

OpenStudy (helder_edwin):

yes. u have to differentiate the numerator as well

OpenStudy (gorica):

or I have to differentiate separately for x and y? I think I have to remind myself on few things :D

OpenStudy (helder_edwin):

the final result would be 1+y' e^y=(1+2yy')/(x+y^2) from this solve for y'

OpenStudy (gorica):

yes, I've got it... But the problem is that a friend of mine that I am doing this for havent's done this yet at faculty and he can't use it. They have just started to do derivatives and I thought this could be solved in an easier way

OpenStudy (helder_edwin):

i got y=(1-x-y^2)/[e^y(x+y^2)-2y]

OpenStudy (helder_edwin):

I see.

OpenStudy (solomonzelman):

and still haven't got the y, when solving for a variable (just reminding, variable should be only on one side. like if you solve for r (in a different question) you can't say r=....r....

OpenStudy (solomonzelman):

I could be wrong, I stink at math, but that's how I was taught in 7th grade.

OpenStudy (blake57roger):

come to the qestoin below please

OpenStudy (blake57roger):

question